OpenStudy (anonymous):

Would greatly appreciate help with this. Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1). Choose one answer. a. parallel: y = -x - 1 perpendicular: y = x + 2 b. parallel: y = x - 1 perpendicular: y = -x + 1 c. parallel: y = x + 1 perpendicular: y = -x - 3 d. parallel: y = 2x - 2 perpendicular: y = -2x - 1

OpenStudy (anonymous):

@Luigi0210 u got this:)

OpenStudy (anonymous):

if no one can help I truly understand! no worries :)

OpenStudy (anonymous):

haha ok ill try to help tho:)

OpenStudy (anonymous):

Only if you are up for it :O

OpenStudy (luigi0210):

Parallel: same slope Perpendicular: negative reciprocal slope

OpenStudy (luigi0210):

So what's the slope of the equation?

OpenStudy (anonymous):

see ur the best @Luigi0210

OpenStudy (luigi0210):

Not even..

OpenStudy (anonymous):

yeahh huh

OpenStudy (anonymous):

isnt in y=mx+b forum the (M) is slope?

OpenStudy (luigi0210):

@proverbsm what is the slope of the original equation?

OpenStudy (luigi0210):

and yes

OpenStudy (anonymous):

oh hi Emily! and I just don't know the proper way to set this up in order to find a solution.

OpenStudy (anonymous):

hi lol:)

OpenStudy (luigi0210):

Once you find the slope, just plug that and the points given into the point-slope formula. y-y1=m(x-x1)

OpenStudy (anonymous):

haha ur so smart @Luigi0210

OpenStudy (anonymous):

Oh thank you very much, I actually forgot that knowledge I once knew! @Luigi0210

OpenStudy (luigi0210):

@emilyhaddad if I was smart I wouldn't of failed art -_- @proverbsm No problem, what will be your final answer?

OpenStudy (anonymous):

@Luigi0210 ok that doesn't mean anything:)

OpenStudy (rmrjr22):

do you get it?

OpenStudy (luigi0210):

@rmrjr22 would of been better for the job than me.

OpenStudy (rmrjr22):

:) thanks, you seem to be doing just fine

OpenStudy (anonymous):

@Luigi0210 you did fine!

OpenStudy (rmrjr22):

think he left.. lol!

OpenStudy (anonymous):

haha yeah:)

OpenStudy (rmrjr22):

@proverbsm u around?

OpenStudy (anonymous):

c. parallel: y = x + 1 perpendicular: y = -x - 3

OpenStudy (anonymous):

whaddup?

OpenStudy (luigi0210):

You are correct, good job

OpenStudy (anonymous):

Thank you!